# Gradient Diffusion: A Perturbation-Resilient Gradient Leakage Attack

### v1

@@ Abstract
Recent years have witnessed the vulnerability of Federated Learning (FL) against gradient leakage attacks, where the private training data can be recovered from the exchanged gradients, making gradient protection a critical issue for the FL training process. Existing solutions often resort to perturbation-based mechanisms, such as differential privacy, where each participating client injects a specific amount of noise into local gradients before aggregating to the server, and the global distribution variation finally conceals the gradient privacy. However, perturbation is not always the panacea for gradient protection since the robustness heavily relies on the injected noise. This intuition raises an interesting question: is it possible to deactivate existing protection mechanisms by removing the perturbation inside the gradients? In this paper, we present the answer: yes and propose the Perturbation-resilient Gradient Leakage Attack (PGLA), the first attempt to recover the perturbed gradients, without additional access to the original model structure or third-party data. Specifically, we leverage the inherent diffusion property of gradient perturbation protection and construct a novel diffusion-based denoising model to implement PGLA. Our insight is that capturing the disturbance level of perturbation during the diffusion reverse process can release the gradient denoising capability, which promotes the diffusion model to generate approximate gradients as the original clean version through adaptive sampling steps. Extensive experiments demonstrate that PGLA effectively recovers the protected gradients and exposes the FL training process to the threat of gradient leakage, achieving the best quality in gradient denoising and data recovery compared to existing models. We hope to arouse public attention on PGLA and its defense.
Xuan Liu* HKPolyU Siqi Cai* WHUT Qihua Zhou SZU
Song Guo HKUST Ruibin Li HKPolyU Kaiwei Lin WHUT
1 Introduction
With the latest success in real-life applications of various data-based generative models, the privacy and information security of these models are under increasing scrutiny[2]. Federated Learning (FL), as a newly proposed privacy preservation distributed machine learning paradigm, collaboratively trains a global model on the parameter server by aggregating computed updates from local clients [18], is widely used in various training scenarios[19, 21, 27], for example, Gboard and Medical Diagnosis System, where data security and privacy are highly significant[30]. Specifically, FL protects users’ privacy by sharing only the model parameters or gradients but not the raw data [15]. Existing FL protocol designs have been demonstrated to be vulnerable to adversaries both inside and outside the system, jeopardizing data privacy and system robustness [17]. Among these, gradient leakage attacks have shown serious privacy hazards [10, 16, 35]. To better protect the private information of the clients from malicious attackers during the training and inference periods, protection of shared gradients becomes an almost essential step [23]. Among existing gradient protection methods, gradient perturbation, as an extensively used privacy-preserving method, has already been frequently applied in FL scenarios efficiently [20] (schematic diagram shown in Fig. 1).
@@ 1 Introduction
Figure 1: Threat model. The FL training process is threatened by gradient leakage attacks, where the attacker can intercept the exchanged gradients $\nabla W$ to recover the private training data. Previous work often protects the gradients by injecting perturbation into the gradients to form $\nabla W^{\prime}_{N}$ and $\nabla W^{Protect}$. Our PGLA removes the perturbation injected in the protected gradients via the diffusion-based denoising model.
@@
Considering that gradient data is the result of linear transformations applied to the original data, such as image data, we posit that if the original data is stable and predictable, then the corresponding gradient data will also be stable and predictable. Moreover, we discovered that the essence of gradient perturbation protection is akin to the diffusion process applied to inherently structurally stable gradient data. Based on these two points, we believe that compared to other models, diffusion models [12] have a natural applicability for gradient perturbation denoising. In this paper, we reveal the natural diffusion properties of gradient perturbation protection and propose PGLA: Perturbation-resilient Gradient Leakage Attack, the first general gradient diffusion attack that is capable of rendering various kinds of perturbations on gradients (e.g. differential privacy [26], certain layer representation perturbation [25], dynamic perturbation [29]) nearly invalid. Based on the structure of obtaining protected shared gradients, our approach involves constructing a surrogate downstream task model for the target attack model. Without access to the original model structure or third-party datasets, we utilize the surrogate gradient data supply model to provide training data for our Gradient Diffusion Model. The trained Gradient Diffusion Model enables us to approximate the original shared gradients of the target clients from the perturbed privacy-preserving gradients obtained by malicious attackers. PGLA embeds the perturbation protection mechanism into the reverse diffusion process and regulates privacy-preserving gradients denoising node as well as the diffusion forward and reverse time steps with the perturbation level $M$ as an adaptable parameter. In summary, the key contributions of our work are:
@@
  • We disclose the natural diffusion process in general gradient perturbation mechanisms by introducing perturbation adaptive parameter $M$.
@@
  • We propose PGLA: Perturbation-resilient Gradient Leakage Attack, which successfully breaks the bottleneck that existing gradient leakage attacks cannot effectively leak privacy under gradient perturbation protection.
  • We present and demonstrate the vulnerability of gradient perturbation protection under diffusion denoising and the suitability of the diffusion model on gradient morphology data.
@@
  • We design the first practical gradient diffusion attack strategy (including surrogate model construction, adjusting gradients to build diffusion training dataset, etc.), and conduct experiments under general perturbation protection FL system. Results show that PGLA achieves the best gradients denoising quality and privacy leakage ability on commonly used image datasets.
2 Preliminary
Federated Learning (FL) provides comparative privacy guarantees for model training which does not require the original data and the participating clients are free to enter or leave [17]. Considering a general FL system with homogeneous architectures. Regarding the gradient protection issue in the FL training process, gradient perturbation is one of the most widely used mechanisms. Here, we present the key preliminary definitions of gradient perturbation and gradient attacks that are pertinent to our work.
Table 1: Explanation of Necessary Notations.
$\nabla W$ Original gradients without any perturbation
$\nabla W^{\prime}$ FL shared gradients protected by perturbation
$\nabla W^{R}$ Recovered gradients of PGLA
$\nabla W_{j}$ Surrogate gradients for PGLA training
$(x_{j},y_{j})$ $Random_{(data,label)}$ for gradients extract
$N$ Number of clients in FL
$F_{i}$ Client’s model (i: $1\rightarrow N$)
$W^{s}$ Surrogate model weight parameter
$F^{s}$ Surrogate model
$X_{0}$ Data input of PGLA training
$\alpha_{t},\beta_{t}$ PGLA forward process variances
$f_{\theta}$ PGLA noise prediction function
$M$ Noise perturbation scale on gradients
$\varepsilon$ DP privacy budget
$\nabla s$ DP sensitivity ($\nabla s_{c}-client$,$\nabla s_{s}-sever$)
$\sigma_{dpc/s}$ DP noise scale (Standard Deviation)
2.1 Differential Privacy (DP)
@@
Generally, DP is defined based on the concept of adjacent database and has been applied in various practical areas in Artificial Intelligence to protect privacy information through adding specific perturbation, e.g. Google’s RAPPORT [8] and large-scale graph data publishing [4]. Tracing back to the definition of Differential Privacy, in this paper we discuss both $\varepsilon-DP$ and $(\varepsilon,\delta)-DP$ [5, 6, 1] as our sample attack background. Here, we present two core definitions first.
@@
Definition 1. A randomized mechanism $M$: $D\rightarrow R$ with domain $D$ and range $R$ satisfies $\varepsilon$ differential privacy if for any two adjacent inputs d, d’ $\in D$ and for any subset of outputs $S\subseteq R$, and it holds that:
$Pr[M(d)\in S]\leq e^{\varepsilon}Pr[M(d^{\prime})\in S].$ (1)
@@
Definition 2. A randomized mechanism $M$: $D\rightarrow R$ with domain $D$ and range $R$ satisfies $(\varepsilon,\delta)$ differential privacy if for any two adjacent inputs d, d’ $\in D$ and for any subset of outputs $S\subseteq R$, and it holds that:
$Pr[M(d)\in S]\leq e^{\varepsilon}Pr[M(d^{\prime})\in S]+\delta.$ (2)
@@
$\delta$-approximation is preferably smaller than $1/|d|$. Note that we usually apply Laplace perturbation for the definition in Eq. (1). However, as to the definition in Eq. (2), the noise perturbation mechanism needs to be Gaussian Mechanisms, which can adapt both $\varepsilon$ and $\delta$. According to Theorem A.1 in Gaussian Mechanisms [7] to ensure the given Gaussian noise distribution $N(0,\sigma^{2}_{dp})$ well preserves $(\varepsilon,\delta)-DP$, the noise scale should satisfy :
$\sigma_{dp}$$\geq$ $c_{dp}\nabla s/\varepsilon,\quad\varepsilon\in(0,1),$ (3)
$c_{dp}$$\geq$ $\sqrt{2ln(1.25/\delta)},$ (4)
where $\nabla s=max_{D,D^{\prime}}||S(D)-S(D^{\prime})||$ and $s$ represents the real-value function. $c_{dp}$ can be considered as a hyperparameter used to define the DP boundary.
2.2 FL with Differential Privacy (FL-DP)
FL-DP protects the shared model parameters or gradients between clients and the server during the FL training process by applying Local Differential Privacy (LDP [33, 26]) and (or) Differential Privacy Stochastic Gradient Descent algorithm (DPSGD [1, 34]).
@@ 2.2 FL with Differential Privacy (FL-DP)
Taking the local participates site for example, the main actions for FL-DP can be divided into four steps: setting DP noise mechanism, local gradient clipping, local gradient perturbation, and uploading the protected parameters to the server [28]. FL involves multiple participants, making composition theorem in DP necessary when applying noise. Among the composition DP methods available, including Simple Composition, Advanced Composition [7], and Moments Accountant [1], we utilize Simple Composition for the following discussion. In the case where $M_{i}$ satisfies $(\varepsilon,\delta)-DP$, the composition ($M_{1},M_{2},\cdots,M_{k}$) satisfies ($\sum_{i=1}^{k}\varepsilon_{i},\sum_{i=1}^{k}\delta_{i}$). Aligning with the state-of-art FL-DP framework [28], We applied Simple Composition with the following noise calculation formulas for local client gradient perturbation, considering both $\varepsilon-DP$ (using the Laplace mechanism) and $(\varepsilon,\delta)-DP$ (using the Gaussian mechanism):
$\textit{Laplace}:\sigma_{dpc}$ $=$ $\nabla s_{c}\times 1/\varepsilon,$ (5)
@@
$\textit{Gaussian}:\sigma_{dpc}$ $=$ $\nabla s_{c}\times\sqrt{2ln(1.25/\delta)}/\varepsilon,$ (6)
@@
where $\nabla s_{c}$ represents the sensitivity and can be formulated as $\nabla s_{c}=\frac{2C}{m}$. Note that $C$ is the clipping threshold for bounding $||\nabla W_{i}||\leq C$, where $\nabla W_{i}$ denotes the unperturbed gradient from the $i-th$ client and $m$ is the minimum size of the local datasets.
From the parameter server site, referring to noising before model aggregation FL (NbAFL) [28], the standard deviation $\sigma_{dps}$ of Gaussian noises that are added to $\nabla W_{server}$ (aggregated gradients at the server to be broadcast to clients) in an $N-client$ FL system is as follows:
$\sigma_{dps}=\left\{\begin{array}[]{cl}\frac{c_{dp}\nabla s_{s}}{\varepsilon}&\quad when\quad T_{dps}>L_{dpc}\sqrt{N}\\ 0&\quad when\quad T_{dps}\leq L_{dpc}\sqrt{N}\\ \end{array}\right.$ (7)
$\nabla s_{s}=\frac{2C}{m}\times\frac{\sqrt{T_{dps}^{2}-L_{dpc}^{2}N}}{N}.$ (8)
The variables $T_{dps}$ and $L_{dpc}$ represent the aggregation times and exposures of uploaded parameters from each client under the FL-DP system, respectively.
2.3 FL with Perturbation Protection (FL-PP)
In addition to DP, FL-PP includes a range of alternative gradient perturbation techniques, which do not need to satisfy DP concepts on similarity and can employ noise to effectively obscure private data. (e.g. adding random global noise to gradients, adding noise to specific layers, and dynamic noise addition) [32].
2.4 Gradient Inversion Attack (GradInv)
@@
$\quad min||\nabla W_{\xi}-\nabla W||:(x_{\xi},y_{\xi})\rightarrow(x^{*}_{\xi},y^{*}_{\xi}).$ (9)
Previous works utilize Peak Signal-to-Noise Ratio (PSNR) of the recovered images to evaluate the performance of GradInv, with the threshold for the success of such attacks typically hinging on the degree of detail discernible to the human eye in the reconstructed private images[22, 9, 35, 16]. To effectively safeguard privacy information against GradInv, a plethora of gradient perturbation methods have been extensively employed [35].
[…]
3 Methodology
PGLA is the first Diffusion Attack Method focusing on the gradient data structure that can be applied to multiple kinds of gradient perturbation protection in FL through adaptive parameters setting on both forward and reverse process of the Gradient Diffusion Model employed in PGLA. The threat models targeted by PGLA encompass various forms of gradient perturbation. Meanwhile, the extremely similar noise mechanism on gradient perturbation protection and diffusion Markovian process provides a mathematical necessity for PGLA to realize an efficient privacy attack on gradient perturbation protection. Overall, our PGLA method (shown in Fig. 2) can be summarized in the following four steps:
(1) Get Protected Gradients. Honest-but-curious malicious attackers steal the Shared Perturbed Gradients $\nabla W^{\prime}$ during the FL training process by hiding on the server side or waiting on the way of parameter sharing (Fig.1).
Algorithm 1 Gradients Extracted for PGLA Training
@@ 3 Methodology
8:  end for
@@
Before delving into Step (3) and Step (4), two configurations that recur in the subsequent steps are defined in advance:
[Gradient Adjustment]: Gradients each with total size $L$ ($L=L_{\nabla W_{Layer1}}+L_{\nabla W_{Layer2}}+...+L_{\nabla W_{LayerN}}$) are adjusted into $1\times g\times g$ ($g=L+P$: $g=$ minimum integer satisfies $g^{2}>L$; $P=$ 0-padding size) before feeding into Gradient Diffusion Model for training or inference to adapt gradient diffusion process. The adjustment is only related to Diffusion procedures, gradients are restored to their original structure and size before Gradient-Based Attacks and Evaluations.
@@
Algorithm 2 Gradient Diffusion Model Training
1:  if With Condition $\nabla W^{\prime}$ then
2:     $X_{0}=(\nabla W^{\prime},\nabla W_{j})$;
3:  else
4:     $X_{0}=\nabla W_{j}$;
5:  end if
6:  Repeat:
7:   $X_{0}\sim p(X_{0})$; $t\sim Uniform(1\rightarrow T)$; $\epsilon\sim N(\textbf{0},\textbf{I})$;
8:   Take a gradient descent step on: $\nabla_{\theta}||\epsilon-f_{\theta}(\sqrt{\gamma_{t}}X_{0}+\sqrt{1-\gamma_{t}}\epsilon,t)||^{2}$;
9:  until converged;
@@
Figure 3: Visualization of Markovian gradient diffusion process. $M$ is the noise scale of perturbation, which is set as the adaptive parameter in [M-Adaptive Process]. PGLA Train and PGLA Inference correspond to Algorithm 2 and Algorithm 3 respectively.
PGLA’s Gradient Diffusion Model is inspired by DDPM[12], to be specific:
In Step (3), refer to Algorithm 2 lines 4 to 7, we set $t\in(1,T)$ as the time steps of Gaussian noise addition. $T$ is the total time step of the forward Markovian diffusion process. $\alpha_{t}(0<\alpha_{t}<1)$ denote the adaptive variables at each iteration and $\gamma_{t}=\prod_{t=0}^{t}\alpha_{t}$. With the above settings, the forward process ($q$: with no learnable parameters) of PGLA’s Gradients Diffusion Model is:
$q(X_{1}|X_{0})$ $=$ $N(X_{1};\sqrt{\alpha_{1}}X_{0},(1-\alpha_{1})I),$ (10)
$q(X_{t}|X_{0})$ $=$ $N(X_{t};\sqrt{\gamma_{t}}X_{0},(1-\gamma_{t})I).$ (11)

### v2

# Mjölnir: Breaking the Shield of Perturbation-Protected Gradients via Adaptive Diffusion

@@ Abstract
Perturbation-based mechanisms, such as differential privacy, mitigate gradient leakage attacks by introducing noise into the gradients, thereby preventing attackers from reconstructing clients’ private data from the leaked gradients. However, can gradient perturbation protection mechanisms truly defend against all gradient leakage attacks? In this paper, we present the first attempt to break the shield of gradient perturbation protection in Federated Learning for the extraction of private information. We focus on common noise distributions, specifically Gaussian and Laplace, and apply our approach to DNN and CNN models. We introduce Mjölnir, a perturbation-resilient gradient leakage attack that is capable of removing perturbations from gradients without requiring additional access to the original model structure or external data. Specifically, we leverage the inherent diffusion properties of gradient perturbation protection to develop a novel diffusion-based gradient denoising model for Mjölnir. By constructing a surrogate client model that captures the structure of perturbed gradients, we obtain crucial gradient data for training the diffusion model. We further utilize the insight that monitoring disturbance levels during the reverse diffusion process can enhance gradient denoising capabilities, allowing Mjölnir to generate gradients that closely approximate the original, unperturbed versions through adaptive sampling steps. Extensive experiments demonstrate that Mjölnir effectively recovers the protected gradients and exposes the Federated Learning process to the threat of gradient leakage, achieving superior performance in gradient denoising and private data recovery.
Introduction
@@
Figure 1: Threat model. The FL training process is threatened by gradient leakage attacks, where the attacker can intercept the exchanged gradients $\nabla W$ to recover the private training data. Previous work often protects the gradients by injecting perturbation into the gradients to form $\nabla W^{\prime}_{N}$ and $\nabla W^{Protect}$. Our Mjölnir removes the perturbation injected in the protected gradients via the adaptive diffusion process.
Federated Learning (FL) is a distributed machine learning paradigm that facilitates collaborative model training without directly transmitting raw training data. Instead, it aggregates gradients or parameters shared among clients to build a global model (McMahan et al. 2017; Gong et al. 2024; Wu et al. 2023; Zhang et al. 2024). This approach preserves the privacy of raw data by keeping it within its originating domain, addressing concerns associated with traditional centralized data processing. However, FL is vulnerable to gradient inversion attacks (Geng et al. 2023; Liu et al. 2023; Zhu, Liu, and Han 2019), in which adversaries can potentially reconstruct sensitive user data from the shared gradients. This vulnerability has spurred significant research into gradient protection techniques (Tan et al. 2024a; Rodríguez-Barroso et al. 2023). Gradient Perturbation, such as differential privacy (DP), injecting noise into gradients to enhance privacy, has been proven to be an effective strategy for safeguarding data in FL scenarios (Wei et al. 2020; Ouadrhiri and Abdelhadi 2022; Wang, Hugh, and Li 2024).
Perturbation-based gradient protection achieves its goal by adding noise to the gradients. If a method is developed to effectively eliminate this noise, the protective mechanism of this approach becomes ineffective. Diffusion Model (Ho, Jain, and Abbeel 2020; Gong 2023) has the natural applicability for denoising perturbation, and is a potential approach to attack perturbation-based gradient protection. We hold the above idea based on two points: 1) Considering that gradient is the result of linear transformations applied to the training data, such as image data, we posit that if the original data is stable and predictable, then the corresponding gradient data will also be stable and predictable. 2) The essence of gradient perturbation protection is akin to the diffusion process applied to inherently structurally stable gradient data.
@@
In this paper, we present the first attempt to break the shield of gradient perturbation protection in FL based on the diffusion model. We reveal the natural diffusion properties of gradient perturbation protection and propose Mjölnir [^1] Mjölnir is the hammer wielded by Thor and is renowned for its incredible power to break any shield., a perturbation-resilient gradient leakage attack method and the first general gradient diffusion attack (schematic diagram shown in Fig. 2) that is capable of rendering various kinds of perturbations on gradients (e.g. Differential Privacy (Truex et al. 2020), certain layer representation perturbation (Sun et al. 2021), dynamic perturbation (Wei et al. 2021)) nearly invalid. As the first attempt, we focus on common noise distributions, specifically Gaussian and Laplace, and apply our approach to DNNs and CNNs. Our method involves constructing a surrogate model for the target attack model to obtain protected shared gradients. In the absence of the original model structure and third-party datasets, we use the surrogate gradient data supply model to generate training data for our gradient diffusion model. This trained gradient diffusion model allows us to approximate the original shared gradients from the perturbed, privacy-preserving gradients obtained by attackers. Mjölnir integrates the perturbation protection mechanism into the reverse diffusion process, adjusting the privacy-preserving gradient denoising nodes and regulating the diffusion’s forward and reverse time steps with the perturbation level $M$ as an adaptable parameter.
In summary, our key contributions include:
@@
  • We disclose the natural diffusion process in general gradient perturbation mechanisms and introduce a perturbation adaptive parameter $M$ to adaptively adjust the diffusion step size according to the degree of perturbation.
@@
  • We propose Mjölnir, the first practical gradient diffusion attack strategy to recover the perturbed gradients, without additional access to the original model structure and third-party data, which breaks the bottleneck that existing gradient leakage attacks cannot effectively leak privacy under gradient perturbation protection.
@@
  • We demonstrate the vulnerability of gradient perturbation protection under the Mjölnir adaptive diffusion denoising process. Experimental results under the general perturbation protection FL system show that Mjölnir achieves the best gradient denoising quality and privacy leakage ability on commonly used image datasets.
Background and Related Work
FL with Perturbation Protection (FL-PP)
FL-PP includes a variety of alternative gradient perturbation techniques, such as adding random global noise to gradients, noise to specific layers, and dynamic noise addition (Zhang et al. 2022). Among these, FL with Differential Privacy (FL-DP) is the most widely used method (Shi et al. 2022; Chen et al. 2023; Tan et al. 2024b).
@@
Generally, DP is defined based on the concept of adjacent database and has been applied in various practical areas in Artificial Intelligence to protect privacy information through adding specific perturbation, e.g. Google’s RAPPORT (Erlingsson, Pihur, and Korolova 2014) and large-scale graph data publishing (Ding et al. 2021). FL-DP protects the shared model parameters or gradients between clients and the server during the FL training process by applying Local Differential Privacy (LDP (Zhao et al. 2021; Truex et al. 2020)) and (or) Differential Privacy Stochastic Gradient Descent algorithm (DPSGD (Abadi et al. 2016; Zhou et al. 2023)). In this paper we discuss both $\varepsilon-DP$ and $(\varepsilon,\delta)-DP$ (Dwork et al. 2006a, b; Abadi et al. 2016) as our sample attack background (Shi et al. 2022; Chen et al. 2023; Tan et al. 2024b):
@@
Assumption 1. A randomized mechanism $M$: $D\rightarrow R$ with domain $D$ and range $R$ satisfies $\varepsilon$ differential privacy if for any two adjacent inputs d, d’ $\in D$ and for any subset of outputs $S\subseteq R$, and it holds that:
$Pr[M(d)\in S]\leq e^{\varepsilon}Pr[M(d^{\prime})\in S]$ (1)
@@
Assumption 2. A randomized mechanism $M$: $D\rightarrow R$ with domain $D$ and range $R$ satisfies $(\varepsilon,\delta)$ differential privacy if for any two adjacent inputs d, d’ $\in D$ and for any subset of outputs $S\subseteq R$, and it holds that:
$Pr[M(d)\in S]\leq e^{\varepsilon}Pr[M(d^{\prime})\in S]+\delta$ (2)
@@
$\delta$-approximation is preferably smaller than $1/|d|$. We usually apply Laplace perturbation for the definition in Eq. (1). However, as to the definition in Eq. (2), the noise perturbation mechanism needs to be Gaussian Mechanisms, which can adapt both $\varepsilon$ and $\delta$. Refer to Theorem A.1 in Gaussian Mechanisms (Dwork and Roth 2014), with $\nabla s=max_{D,D^{\prime}}||S(D)-S(D^{\prime})||$ ($s$ is the real-value function) and $c_{dp}$ denotes the hyperparameter used to define the DP boundary, to ensure Gaussian noise distribution $N(0,\sigma^{2}_{dp})$ well preserves $(\varepsilon,\delta)-DP$, the noise scale should satisfy:
$\sigma_{dp}$$\geq$ $c_{dp}\nabla s/\varepsilon,\quad\varepsilon\in(0,1)$ (3)
$c_{dp}$$\geq$ $\sqrt{2ln(1.25/\delta)}$ (4)
@@
Taking the local participates site for example, the main actions for FL-DP can be divided into four steps: setting DP noise mechanism, local gradient clipping, local gradient perturbation, and uploading the protected parameters to the server (Wei et al. 2020). FL involves multiple participants, making composition theorem in DP necessary when applying noise. Among the composition DP methods available, including Simple Composition, Advanced Composition (Dwork and Roth 2014), and Moments Accountant (Abadi et al. 2016), we utilize Simple Composition for the following discussion. In the case where $M_{i}$ satisfies $(\varepsilon,\delta)-DP$, the composition ($M_{1},M_{2},\cdots,M_{k}$) satisfies ($\sum_{i=1}^{k}\varepsilon_{i},\sum_{i=1}^{k}\delta_{i}$). Aligning with the state-of-art FL-DP framework (Wei et al. 2020; Dwork et al. 2006a, b; Abadi et al. 2016; Shi et al. 2022; Chen et al. 2023; Tan et al. 2024b), We applied the following assumptions for noise calculation in local client gradient perturbation, considering both $\varepsilon$-DP (using the Laplace mechanism) and $(\varepsilon,\delta)$-DP (using the Gaussian mechanism):
$\textit{Laplace}:\sigma_{dpc}$ $=$ $\nabla s_{c}\times 1/\varepsilon$ (5)
@@
$\textit{Gaussian}:\sigma_{dpc}$ $=$ $\nabla s_{c}\times\sqrt{2ln(1.25/\delta)}/\varepsilon$ (6)
@@
where $\nabla s_{c}$ is the sensitivity and can be formulated as $\nabla s_{c}=\frac{2C}{m}$. $C$ is the clipping threshold for bounding $||\nabla W_{i}||\leq C$, where $\nabla W_{i}$ denotes the unperturbed gradient from the $i-th$ client and $m$ denotes the minimum size of local datasets.
Gradient Inversion Attack (GradInv)
@@
$\quad min||\nabla W_{\xi}-\nabla W||:(x_{\xi},y_{\xi})\rightarrow(x^{*}_{\xi},y^{*}_{\xi})$ (7)
Previous works utilize Peak Signal-to-Noise Ratio (PSNR) of the recovered images to evaluate the performance of GradInv, with the threshold for the success of such attacks typically hinging on the degree of detail discernible to the human eye in the reconstructed private images (Geiping et al. 2020; Zhu, Liu, and Han 2019; Liu et al. 2023).
Methodology
Figure 2: Mjölnir Overview. After intercepting the exchange gradients protected by unknown perturbation ($\nabla W^{\prime}$) from clients, the attacker will (1) leverage the invariance of gradient data structure before and after perturbation to construct the surrogate model; (2) feed random image dataset into the surrogate model to extract surrogate clean gradients; (3) flatten the surrogate gradients and pad them to the appropriate size $g^{2}=P+L$ (g is the minimum integer satisfies $g^{2}>L$) to create a surrogate gradient ($\nabla W^{s}$) dataset for training Gradient Diffusion Model (if additional conditions are chosen to guide the training process, joint ($\nabla W^{s}$,$\nabla W^{s}_{perturbed}$) or ($\nabla W^{s}$,$\nabla W^{\prime}$) as the training dataset, where $\nabla W^{s}_{perturbed}$ denote surrogate gradients applied with known perturbation; if no conditions are needed, directly use $\nabla W^{s}$); (4) use the trained Gradient Diffusion Model to denoise $\nabla W^{\prime}$ to generate the recovered gradient $\nabla W^{R}$. From $\nabla W^{R}$, the attacker can recover clients’ privacy information.
Mjölnir is the first Diffusion Attack Method focusing on the gradient data structure that can be applied to multiple kinds of gradient perturbation protection in FL through adaptive parameters setting on both forward and reverse process of the Gradient Diffusion Model employed in Mjölnir. The threat models targeted by Mjölnir encompass various forms of gradient perturbation. Meanwhile, the extremely similar noise mechanism on gradient perturbation protection and diffusion Markovian process provides a mathematical necessity for Mjölnir to realize an efficient privacy attack on gradient perturbation protection. Overall, our Mjölnir method (shown in Fig. 2) can be summarized in four steps:
Step (1) Get Protected Gradients. Honest-but-curious malicious attackers steal the Shared Perturbed Gradients $\nabla W^{\prime}$ during the FL training process by hiding on the server side or waiting on the way of parameter sharing (Fig.1).
Algorithm 1 Gradients Extracted for Mjölnir Training
@@
Step (2) Construct Surrogate Model. Construct Surrogate Gradients Data Supply Model $F^{s}$ from the data structure of Shared Perturbed Gradients $\nabla W^{\prime}$ stolen by the attacker that can output the same gradient data structure as the target attacked client’s local model (Algorithm 1, Fig. 2).
Before delving into Step (3) and Step (4), two configurations that recur in the subsequent steps are defined:
[Gradient Adjustment]: Gradients each with total size $L$ ($L=L_{\nabla W_{Layer1}}+L_{\nabla W_{Layer2}}+...+L_{\nabla W_{LayerN}}$) are adjusted into $1\times g\times g$ ($g=L+P$: $g=$ minimum integer satisfies $g^{2}>L$; $P=$ 0-padding size) before feeding into Gradient Diffusion Model for training or inference to adapt gradient diffusion process. The adjustment is only related to Diffusion procedures, gradients are restored to their original structure and size before further Gradient-Based Attacks and Evaluations.
Algorithm 2 Gradient Diffusion Model Training
1:if With Condition $\nabla W^{\prime}$ then $X_{0}=(\nabla W^{\prime},\nabla W_{j})$;
@@
3:Repeat:
4:$X_{0}\sim p(X_{0})$; $t\sim Uniform(1\rightarrow T)$; $\epsilon\sim N(\textbf{0},\textbf{I})$;
5: Take a gradient descent step on: $\nabla_{\theta}||\epsilon-f_{\theta}(\sqrt{\gamma_{t}}X_{0}+\sqrt{1-\gamma_{t}}\epsilon,t)||^{2}$;
6:until converged;
Step (3) Train Gradients Diffusion Model. Construct our Gradient Diffusion Model that takes into account the level of knowledge about the attacked model (e.g., whether the level of perturbation noise or the type of noise is known). Conduct [Gradient Adjustment] to Clean Gradients $\nabla W^{s}$ extracted from Surrogate Model $F^{s}$ to build a training gradient dataset to train our Gradient Diffusion Model (Fig. 2, Fig.3). Algorithm 2 demonstrates a detailed training process.
Step (4) Recover Original Gradients. The stolen Shared Perturbed Gradients $\nabla W^{\prime}$ are put into our trained Mjölnir Gradient Diffusion Model after [Gradient Adjustment] and [M-Adaptive Process] to generate the Recovered Gradients $\nabla W^{R}$ for further Gradient-Based Attack to get certain privacy information based on various demands (Algorithm 3, Fig. 2 and Fig. 3).
Figure 3: Visualization of Markovian gradient diffusion process. $M$ is the noise scale of perturbation, which is set as the adaptive parameter in [M-Adaptive Process]. Mjölnir Train and Mjölnir Inference correspond to Algorithm 2 and Algorithm 3 respectively.
Algorithm 3 Generate Original Gradient
1:if Known Noise Scale $M$ then $\textbf{c}=\frac{1}{\sqrt{1+M^{2}}}$;
@@
3:$X_{T^{\prime}}\sim\textbf{c}\nabla W^{\prime}$;
4:for $t=T^{\prime},...,1$ do
5:     if $t>1$ then $\textbf{z}\sim N(\textbf{0},\textbf{I})$;
@@
7:     $X_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}(X_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\gamma_{t}}}f_{\theta}(X_{t},t))+\sqrt{1-\alpha_{t}}\textbf{z}$;
8:return $X_{0}$; $\nabla W^{R}\leftarrow X_{0}$: $\triangleright$ For Further GradInv;
Mjölnir’s Gradient Diffusion Model is inspired by DDPM(Ho, Jain, and Abbeel 2020), to be specific:
In Step (3) Train Gradients Diffusion Model, refer to Algorithm 2 lines 4 to 6, we set $t\in(1,T)$ as the time steps of Gaussian noise addition. $T$ is the total time step of the forward Markovian diffusion process. $\alpha_{t}(0<\alpha_{t}<1)$ denote the adaptive variables at each iteration and $\gamma_{t}=\prod_{t=0}^{t}\alpha_{t}$. With the above settings, the forward process ($q$: with no learnable parameters) of Mjölnir’s Gradients Diffusion Model can be modeled as:
$q(X_{1}|X_{0})$ $=$ $N(X_{1};\sqrt{\alpha_{1}}X_{0},(1-\alpha_{1})I)$ (8)
$q(X_{t}|X_{0})$ $=$ $N(X_{t};\sqrt{\gamma_{t}}X_{0},(1-\gamma_{t})I)$ (9)